# Detection of evolutionary shifts in variance under an Ornstein–Uhlenbeck model

### v1

@@
Detection of evolutionary shifts in variance under an Ornstein–Uhlenbeck model
[…]
${}^{1,2,3}$epartment of Mathematics and Statistics, Dalhousie University, Nova Scotia, Canada
$*{}^{*}$hese authors have equal contribution
@@
  1. 1.
    Abrupt environmental changes can lead to evolutionary shifts in not only mean (optimal value), but also variance of descendants in trait evolution. There are some methods to detect shifts in optimal value but few studies consider shifts in variance.
  2. 2.
    We use a multi-optima and multi-variance OU process model (Beaulieu, Jhwueng, Boettiger, and O’Meara, 2012) to describe the trait evolution process with shifts in both optimal value and variance and provide analysis of how the covariance between species changes when shifts in variance occur along the path.
  3. 3.
    We propose a new method to detect the shifts in both variance and optimal values based on minimizing the loss function with $\ell_{1}$ penalty. We implement our method in a new R package, ShiVa (Detection of evolutionary shifts in variance).
  4. 4.
    We conduct simulations to compare our method with the two methods considering only shifts in optimal values ($\ell$1ou; PhylogeneticEM). Our method shows strength in predictive ability and includes far fewer false positive shifts in optimal value compared to other methods when shifts in variance actually exist. When there are only shifts in optimal value, our method performs similarly to other methods. We applied our method to the cordylid data ( Broeckhoven, Diedericks, Hui, Makhubo, and Mouton, 2016). ShiVa outperformed $\ell$1ou and phyloEM, exhibiting the highest log-likelihood and lowest BIC.
Keywords— evolutionary shift detection, Ornstein-Uhlenbeck model, LASSO, trait evolution, ensemble method, phylogenetic comparative methods
@@ 1 Introduction
The trait evolution process is related to the environment. Abrupt environmental changes for a species might lead to changes in the evolutionary process of the descendants to adapt to the environment (Losos, 2011; Mahler, Ingram, Revell, and Losos, 2013). The changes in the evolutionary process will be reflected in the observed trait values of present-day species. By detecting abrupt changes based on observed traits, we can get knowledge of unobserved historical environmental changes and better understand the evolutionary history. The most popular statistical models to model the evolutionary process along a phylogenetic tree are Brownian Motion (BM, Felsenstein, 1985) and the Ornstein-Uhlenbeck process (Hansen, 1997). Abrupt changes in the evolutionary process are mainly modeled by shifts in parameters in these models. Most research in this area focuses on the detection of shifts in mean (BM model) or optimal value (OU model). Butler and King (2004) formulate the multiple-optima OU model for adaptive evolution. They assume that the optimum is constant along a branch and may differ between branches due to discrete events where changes in selective regime occur. They test the difference in optima between branches using hypothesis testing. Uyeda and Harmon (2014) propose a Bayesian framework to detect shifts in the selective optimum of OU models. Ho and Ané (2014) illustrate the limitation of AIC and BIC in the shift detection task and introduce modified BIC (Zhang and Siegmund, 2006) into the shift detection task. They perform forward-backward selection to conduct model selection. Khabbazian et al. (2016) formulate the shift detection problem into a variable selection problem and combine the OU model with LASSO to detect the shift points, which they implement in the $\ell$1ou R package. The application of LASSO makes it a fast method. Further, they propose a new information criterion pBIC with correction for phylogenetic correlations. Bastide et al. (2017) develop a maximum likelihood estimation procedure based on the EM algorithm, which they implement in the R package PhylogeneticEM. Zhang et al. (2022) propose an ensemble variable selection method (R package ELPASO). They conduct simulations for scenarios when the common model assumptions are violated and show that measurement error, tree construction error and shifts in the diffusion variance impact the performances of the current methods.
However, these methods only consider shifts in optimal values (mean) without considering shifts in variances. In reality, when abrupt environmental shifts occur, they are likely to not only affect the optimal values but also affect the variances of traits. There are some existing models which allow variance ($\sigma^{2}$) to change. Beaulieu et al. (2012) generalize the Hansen model to allow multiple variance parameters and multiple attraction parameters for different selective regimes. Mitov et al. (2020) explore a sub-family of Gaussian phylogenetic models, denoted $\mathcal{G}_{LInv}$, which support models with shifts in variance and propose a fast likelihood calculation algorithm for the family. Bastide et al. (2021) propose a Bayesian inference framework under a Gaussian evolution model that exploits Hamiltonian Monte Carlo (HMC) sampling. The framework allows a variety of models including models with multiple variance parameters. The above models allow changes in variance parameter, but do not incorporate advanced variable selection methods for determining where these shifts in variance occur. In this paper, we propose a method that can simultaneously detect both shifts in optimal value and shifts in variance.
We use a multi-optima and multi-variance OU process model (Beaulieu, Jhwueng, Boettiger, and O’Meara, 2012) to describe the trait evolution process with shifts in both optimal value and variance. We provide analysis of how the covariance between species changes when shifts in variance occur along the path. We then propose a new method to detect both shifts in optimal value and shifts in variance using LASSO. We implement our method in a new R pacakge, ShiVa (Detection of evolutionary shifts in variance). We conduct simulations and compare with $\ell$1ou and PhylogeneticEM. The remainder of the paper is organized as follows. Section 2 introduces the trait evolution model considering shifts in both optimal value and variance, and formulates the shift detection problem. Section 3 describes our algorithm for shift detection. Section 4 presents our simulation results. The methods are compared using both true positive versus false positive curves of shifts in optimal values and predictive log-likelihood. Section 5 illustrates the use of ShiVa with cordylid data (Broeckhoven, Diedericks, Hui, Makhubo, and Mouton, 2016). Section 6 provides the conclusion and discussion.
@@ 2.1 Trait evolution models
The phylogenetic tree is reconstructed from DNA sequences and assumed known in this paper. The phylogenetic tree reveals the correlation structure between trait values of different species. The trait values are correlated based on the shared evolutionary history of species. The trait values of internal nodes are hidden and only trait values of tip nodes can be observed.
@@ 2.2 Evolutionary shifts in optimal value and variance
For model and computational simplicity, we assume that $\sigma^{2}$ is constant over a branch like the assumption for $\theta$. That is, we assume that any shift in variance on a branch occurs at the beginning of that branch. For shifts in mean, this assumption did not limit the space of possible models due to the unidentifiability. However, for shifts in variance, this does restrict the space of models for the trait values at the leaves, and could therefore adversely affect shift detection. We will study whether this adverse effect is cause for concern in our simulation studies. Under our assumptions, we let $\sigma^{2}_{b}$ denote the variance on the branch $b$, $\Delta\sigma^{2}_{b}$ denote the magnitude of shift in variance on the branch $b$: if no shift occurs on the branch $b$, $\Delta\sigma^{2}_{b}=0$. The covariance between species $i$ and $j$ is given by:
@@
It consists of 3 terms. The first term is the original covariance without any shift. The last 2 terms show the influence of shifts in variances. In addition to the shift sizes $\gamma_{b}$, the second term is influenced by the phylogentic structure ($\mathbf{V}_{i,j}$), higher original covariance between two species leads to larger change in covariance for a fixed shift size (Figure 2); the third term is influenced by the start time of branch b ($\mathbf{q}_{b}$), earlier shifts lead to larger change in covariance between two species (Figure 2).
@@ 3.1 Optimization
We assume that $\alpha$ is known in this paper. The objective here is to find $\mathbf{\beta}$, $\mathbf{\gamma}$ and $\sigma_{0}^{2}$ that minimize Equation 3. When $\mathbf{\gamma}$ is fixed, it can be transformed into a basic LASSO problem. For the optimization of $\mathbf{\gamma}$, we adopt coordinate proximal gradient descent (Parikh, 2014). The proximal gradient algorithm is commonly used to solve non-differentiable convex optimization problems. The objective function is usually broken into two parts: a differentiable part and a non-differentiable part to apply the proximal gradient algorithm. For the optimization, a proximal operator is used to reduce the non-differentiable part.
@@
$L(\gamma_{k})=g(\gamma_{k})+\lambda_{2}\|\gamma_{k}\|$
@@
$\qquad+\frac{1}{2}q_{k}\mathbf{X_{k}^{T}\Sigma^{-1}X_{k}}$
@@
$\text{prox}_{\lambda_{2}h}(z)=S_{\lambda_{2}}(z)=\begin{cases}z-\lambda_{2}&z>\lambda_{2}\\ 0&-\lambda\leq z\leq\lambda_{2}\\ z+\lambda_{2}&z<-\lambda\end{cases}$
The proximal algorithm here steps in the direction $\nabla g(\gamma)$ but setting any values that are close to zero equal to zero. For $\sigma^{2}_{0}$, there is no penalty on the parameter. Therefore, we use gradient descent to update $\sigma_{0}^{2}$. The gradient of $\sigma_{0}^{2}$ is given by:
$\frac{\partial L}{\partial\sigma_{0}^{2}}$ $=-\frac{1}{2}\mathbf{r^{T}}\mathbf{\Sigma}^{-1}\mathbf{R}\mathbf{\Sigma}^{-1}\mathbf{r}+\frac{1}{2}\text{tr}\left(\mathbf{R}\mathbf{\Sigma}^{-1}\right)$
We initialize $\beta$ and $\gamma$ as 0. In each iteration, we first update the covariance matrix $\Sigma$ and transform $\mathbf{X}$ and $\mathbf{Y}$ with the current $\Sigma^{-1/2}$. Then we update $\beta$ by applying LASSO on the transformed data. Then we update $\gamma$ using the proximal algorithm in an elementwise manner. After that, we update $\sigma_{0}^{2}$ using gradient descent. The algorithm is summarized in Algorithm 1.
Algorithm 1 Optimization with $\beta$, $\gamma$ and $\sigma_{0}^{2}$
@@
2:      $M$: maximum number of steps; $t$: step size; $\epsilon$: error tolerance; $\mathbf{V}$; $\mathbf{R}$; $\mathbf{Y}$; $\mathbf{X}$
@@
4:      $\beta_{i}^{0}\leftarrow 0$, $i=1,\ldots,p$ $\gamma_{i}^{0}\leftarrow 0$, $i=1,\ldots,p$ $\sigma_{0}^{2}\leftarrow 1$
@@
6:    $\mathbf{\Sigma}\leftarrow\sigma_{0}^{2}\mathbf{R}+\left(\sum_{i=1}^{p}\gamma_{i}\mathbf{X_{i}X_{i}^{T}}\right)\odot\mathbf{V}+\sum_{i=1}^{p}\gamma_{i}q_{i}\mathbf{X_{i}X_{i}^{T}}$
@@
14:    calculate the gradient for $\sigma_{0}^{2}$: $\frac{\partial L}{\partial\sigma_{0}^{2}}$
15:    update $\sigma_{0}^{2}\leftarrow\sigma_{0}^{2}-t\frac{\partial L}{\partial\sigma_{0}^{2}}$
16:    if max update for $\beta$, $\gamma$ and $\sigma_{0}^{2}$ in the iteration $<\epsilon$ then
17:       break
18:    end if
19:end for
@@ 3.2 Model selection
We have two tuning parameters ($\lambda_{1}$ and $\lambda_{2}$). Because of the efficiency of computing the LASSO path in regression models, we can efficiently tune $\lambda_{1}$ by cross-validation. So for each fixed $\lambda_{2}$ value, we set $\lambda_{1}$ to 0 to obtain an estimation of the covariance matrix. Then we tune $\lambda_{1}$ by cross validation according to the estimated coariance matrix. We then use BIC to select the best value of $\lambda_{2}$.

### v2

@@
Abstract
Abrupt environmental changes can lead to evolutionary shifts in not only the optimal trait value ($\theta$), but also the rate of adaptation ($\alpha$) and the diffusion variance ($\sigma^{2}$) in trait evolution. While several methods exist for detecting shifts in optimal values, few explicitly model shifts in both evolutionary variance and adaptation rates. We use a multi-optima and multi-variance Ornstein-Uhlenbeck (OU) process model (Beaulieu, Jhwueng, Boettiger, and O’Meara, 2012) to describe trait evolution with shifts in both $\theta$ and $\sigma^{2}$ and analyze how covariance between species is affected when shifts in variance occur along the phylogeny. We propose a new method that simultaneously detects shifts in both variance and optimal values by formulating the problem as a variable selection task using an $\ell_{1}$-penalized loss function. Our method is implemented in the R package ShiVa (Detection of evolutionary shifts in variance). Through simulations, we compare ShiVa with methods that only consider shifts in optimal values ($\ell$1ou; PhylogeneticEM). Our method demonstrates improved predictive ability and significantly reduces false positives in detecting optimal value shifts when variance shifts are present. When only shifts in optimal value occur, our method performs comparably to existing approaches. Applying ShiVa to empirical data from cordylid lizards (Broeckhoven, Diedericks, Hui, Makhubo, and Mouton, 2016a), we find that it outperforms $\ell$1ou and PhylogeneticEM, achieving the highest log-likelihood and lowest BIC. evolutionary shift detection, Ornstein-Uhlenbeck model, LASSO, trait evolution, phylogenetic comparative methods
1,2,3Department of Mathematics and Statistics, Dalhousie University, Nova Scotia, Canada
These authors have equal contribution
@@ 1 Introduction
The process of trait evolution is influenced by environmental conditions. When species experience abrupt environmental changes, their evolutionary processes may also change to adapt to the new environment (Losos, 2011; Mahler, Ingram, Revell, and Losos, 2013a). These changes in the evolutionary process leave signatures in the observed traits of present-day species. By detecting such evolutionary shifts based on observed trait data, we gain valuable insights into historical environmental changes and the evolutionary history of species.
To model the evolutionary process along a phylogenetic tree, researchers commonly use two statistical frameworks: the Brownian Motion (BM) model (Felsenstein, 1985) and the Ornstein-Uhlenbeck (OU) process model (Hansen, 1997). Evolutionary shifts are typically modeled as changes in the parameters of these processes. Much of the existing research focuses on detecting shifts in the current value (under BM) or shifts in the optimal trait value (under OU). Butler and King (2004) formulated the multiple-optima OU model, which allows shifts in the selective optima across different branches of a phylogeny and detects these shifts using hypothesis testing. Uyeda and Harmon (2014) introduced a Bayesian framework to detect such shifts in selective optima. Ho and Ané (2014) highlighted limitations of using AIC and BIC in shift detection tasks, proposing a modified BIC and a forward-backward model selection procedure. Khabbazian et al. (2016) treated the shift detection problem as a variable selection task, applying LASSO to detect shifts in selective optima, implemented in the $\ell$1ou R package. This LASSO-based approach enables efficient shift detection, and they further proposed a new information criterion (pBIC) to account for phylogenetic correlations. Other methods, such as PhylogeneticEM (Bastide, Mariadassou, and Robin, 2017), apply expectation-maximization (EM) to estimate shifts in the optima. Zhang et al. (2024) extended this line of work with the ELPASO package, which uses ensemble variable selection to improve robustness under model violations.
However, these methods primarily focus on detecting shifts in optimal values and do not explicitly detect the shifts in evolutionary diffusion variance. In reality, abrupt environmental changes are likely to influence not only the optimal trait values but also the evolutionary variance, which reflects how much random fluctuation occurs around the evolutionary process. In this paper, the term variance specifically refers to the diffusion variance ($\sigma^{2}$), which quantifies the background rate of stochastic variation in trait evolution. This variance may increase under unstable environments or decrease under stabilizing conditions. Some existing methods allow modeling shifts in evolutionary parameters, including the optimal value ($\theta$), the rate of adaptation ($\alpha$) and the evolutionary variance ($\sigma^{2}$). Pagel and Meade (2006) introduced BayesTraits, which allows shifts in both optimal values and rates of evolution (modeled as $\alpha$). Beaulieu et al. (2012) developed the OUwie package, which supports detecting shifts in both optimal values and the rate of adaptation across branches in the tree. Later, Clavel et al. (2015) introduced mvMorph, which extends evolutionary modeling to multivariate traits, allowing shifts in both the rate of adaptation and evolutionary variance across lineages. More recently, Bastide et al. (2021) developed a flexible Bayesian framework using Hamiltonian Monte Carlo (HMC) sampling, which supports models with shifting evolutionary variances. Finally, Mitov et al. (2020) proposed the $\mathcal{G}_{LInv}$ family of Gaussian phylogenetic models, providing an efficient likelihood-based approach to model variance shifts across branches.
While these tools allow evolutionary shifts to be modeled, they typically require either pre-specified regimes or rely on computationally intensive full Bayesian inference. Furthermore, few methods explicitly focus on detecting both shifts in optimal value and shifts in variance simultaneously using a data-driven variable selection framework. In this paper, we propose a novel LASSO-based method that detects both types of shifts within a unified variable selection framework. Our method uses a multi-optima and multi-variance OU process model, similar to the models used in OUwie (Beaulieu, Jhwueng, Boettiger, and O’Meara, 2012), but innovates by formulating the detection of both shifts in optima and shifts in variance as a single variable selection problem, solvable using LASSO. This approach allows flexible and efficient detection of shifts across the entire tree, without requiring a priori specification of regime boundaries.
We implement this approach in a new R package, ShiVa (Detection of Evolutionary Shifts in Variance). Through simulation studies, we compare ShiVa to existing methods, including $\ell$1ou and PhylogeneticEM, evaluating performance in terms of both shift detection accuracy and predictive performance.
The remainder of the paper is organized as follows. The Trait evolution with shifts in both optimal value and variance section introduces the trait evolution model incorporating shifts in both optimal value and variance, and formally defines the shift detection problem. The Methods section describes the LASSO-based shift detection algorithm. The Simulations section presents simulation results, including comparisons with existing methods. The Case study section illustrates the application of ShiVa to empirical data from cordylid lizards (Broeckhoven, Diedericks, Hui, Makhubo, and Mouton, 2016a). Conclusion concludes with a discussion of the advantages, limitations, and potential extensions of the method.
@@ 2.1 Trait evolution models
The phylogenetic tree is reconstructed from DNA sequences and is assumed to be known in this paper. The phylogenetic tree reveals the correlation structure between trait values of different species. The trait values are correlated based on the shared evolutionary history of species. The trait values of internal nodes are hidden and only trait values of tip nodes can be observed. When $\alpha$ is unknown, a reasonable estimate can be obtained by fitting a null model (a model without shifts) using the phylolm package. Throughout this paper, we focus on estimating the parameters related to shifts, while the trait values of internal nodes remain hidden and only the trait values at the tips of the tree are observed.
@@
Brownian motion (BM) was first applied to model the evolution of continuous traits over time by Felsenstein (1985). Under this model, the trait value evolves following a BM process. The process can be written as a stochastic differential equation:
$dX_{t}=\sigma(t)\,dW_{t},$
@@ 2.2 Evolutionary shifts in optimal value and variance
For model and computational simplicity, we assume that $\sigma^{2}$ is constant over a branch like the assumption for $\theta$. That is, we assume that any shift in variance on a branch occurs at the beginning of that branch. For shifts in mean, this assumption did not limit the space of possible models due to the unidentifiability. However, for shifts in variance, this does restrict the space of models for the trait values at the leaves, and could therefore adversely affect shift detection. We will study whether this adverse effect is cause for concern in our simulation studies.
Under our assumptions, we let $\sigma^{2}_{b}$ denote the variance on the branch $b$, $\Delta\sigma^{2}_{b}$ denote the magnitude of shift in variance on the branch $b$: if no shift occurs on the branch $b$, $\Delta\sigma^{2}_{b}=0$. The covariance between species $i$ and $j$ is given by:
@@
It consists of 3 terms. The first term is the original covariance without any shift. The last 2 terms show the influence of shifts in variances. In addition to the shift sizes $\gamma_{b}$, the second term is influenced by the phylogentic structure ($\mathbf{V}_{i,j}$), higher original covariance between two species leads to larger change in covariance for a fixed shift size (Fig. 2); the third term is influenced by the start time of branch b ($\mathbf{q}_{b}$), earlier shifts lead to larger change in covariance between two species (Fig. 2).
@@ 3.1 Optimization
In this paper, we do not estimate the parameter $\alpha$ jointly with other parameters, but instead treat it as fixed using an ad-hoc estimate obtained from the null model fitted by the phylolm R package. Therefore, in this section, we treat it as fixed during the model fitting process. In the simulations, we demonstrate that the estimation of $\alpha$ does not significantly affect the detection results. Therefore, we omit the estimation of $\alpha$ in the ShiVa procedure. In cases where $\alpha$ is unknown, we use an ad-hoc estimate of $\alpha$ provided by fitting a null model (a model without shifts) using the phylolm package. This estimated $\alpha$ is then fixed when running ShiVa. The primary objective here is to find the parameters $\mathbf{\beta}$, $\mathbf{\gamma}$, and $\sigma_{0}^{2}$ that minimize the loss function described in Equation 3. When $\mathbf{\gamma}$ is fixed, the problem reduces to a standard LASSO formulation, which can be efficiently solved.
For optimizing $\mathbf{\gamma}$, we employ coordinate-wise proximal gradient descent, as described by Parikh and Boyd (2014). The proximal gradient algorithm is a powerful tool for handling non-differentiable convex optimization problems. It decomposes the objective function into two components: a differentiable part and a non-differentiable part, allowing for iterative updates that combine gradient descent with the proximal operator. The proximal operator serves to handle the non-differentiable component, promoting sparsity and effectively reducing the influence of irregularities in the model. By applying the proximal gradient algorithm in our optimization process, we can efficiently minimize the objective function even when the problem involves non-smooth terms, such as LASSO regularization.
@@
$L(\gamma_{k})=g(\gamma_{k})+\lambda_{2}\|\gamma_{k}\|_{1}.$
@@
$\qquad+\frac{1}{2}q_{k}\mathbf{X_{k}^{T}\Sigma^{-1}X_{k}},$
@@
$\text{prox}_{\lambda_{2}h}(z)=S_{\lambda_{2}}(z)=\begin{cases}z-\lambda_{2}&z>\lambda_{2},\\ 0&-\lambda_{2}\leq z\leq\lambda_{2},\\ z+\lambda_{2}&z<-\lambda_{2}.\end{cases}$
The proximal algorithm here steps in the direction $\nabla g(\gamma)$ but setting any values that are close to zero equal to zero. For $\sigma^{2}_{0}$, there is no penalty on the parameter. Therefore, we use gradient descent to update $\sigma_{0}^{2}$. To avoid negative values of $\sigma_{0}^{2}$, we use $\tau_{0}=\log(\sigma_{0}^{2})$ for optimization. The gradient of $\tau_{0}$ is given by:
$\frac{\partial L}{\partial\tau}$ $=\left(-\frac{1}{2}\mathbf{r^{T}}\mathbf{\Sigma}^{-1}\mathbf{R}\mathbf{\Sigma}^{-1}\mathbf{r}+\frac{1}{2}\text{tr}\left(\mathbf{R}\mathbf{\Sigma}^{-1}\right)\right)e^{\tau_{0}}.$
We initialize $\beta$ and $\gamma$ as 0. In each iteration, we first update the covariance matrix $\Sigma$ and transform $\mathbf{X}$ and $\mathbf{Y}$ with the current $\Sigma^{-1/2}$. Then we update $\beta$ by applying LASSO on the transformed data. Then we update $\gamma$ using the proximal algorithm in an elementwise manner. After that, we update $\tau_{0}^{2}$ using gradient descent. The algorithm is summarized in Algorithm 1.
Algorithm 1 Optimization with $\beta$, $\gamma$ and $\tau_{0}$
@@
2:      $M$: maximum number of steps; $t$: step size; $\epsilon$: error tolerance; $\mathbf{V}$; $\mathbf{R}$; $\mathbf{Y}$; $\mathbf{X}$; $\lambda_{2}$
@@
4:      $\beta_{i}^{0}\leftarrow 0$, $i=1,\ldots,p$ $\gamma_{i}^{0}\leftarrow 0$, $i=1,\ldots,p$ $\tau_{0}\leftarrow 0$ $L\leftarrow Inf$
@@
6:    $\mathbf{\Sigma}\leftarrow e^{(\tau_{0})}\mathbf{R}+\left(\sum_{i=1}^{p}\gamma_{i}\mathbf{X_{i}X_{i}^{T}}\right)\odot\mathbf{V}+\sum_{i=1}^{p}\gamma_{i}q_{i}\mathbf{X_{i}X_{i}^{T}}$
@@
14:    calculate the gradient for $\tau_{0}$: $\frac{\partial L}{\partial\tau_{0}}$
15:    update $\tau_{0}\leftarrow\tau_{0}-t\frac{\partial L}{\partial\tau_{0}}$
16:    update the loss function $L$ with Equation 3
17:    if update for the loss function $L$ in the iteration $<\epsilon$ then
18:       break
19:    end if
20:end for
@@ 3.2 Model selection
For the model selection process, we use a strategy to fine-tune the parameters $\lambda_{1}$ and $\lambda_{2}$ to strike an optimal balance between model simplicity and performance.
  • Fixing $\lambda_{2}$: We begin by setting $\lambda_{2}$ to a fixed value, which controls the degree of regularization for variance shifts. With $\lambda_{2}$ fixed, we focus on adjusting $\lambda_{1}$, which regulates the penalization of shifts in the optimal values.
  • Cross-validation for $\lambda_{1}$: For each fixed value of $\lambda_{2}$, we use cross-validation to find the best $\lambda_{1}$. We use the cv.glmnet function from the R package glmnet to efficiently perform cross-validation here. We use cross-validation because cv.glmnet allows for quick and easy computation of cross-validation results, making it more convenient and efficient for selecting the optimal parameters.
  • Iterating over different $\lambda_{2}$ values: After optimizing $\lambda_{1}$ for one fixed $\lambda_{2}$, we repeat this process for multiple values of $\lambda_{2}$, producing a set of candidate models, each associated with different pairs of $\lambda_{1}$ and $\lambda_{2}$.
  • Selecting the best model using BIC: To identify the optimal model, we compare all candidate models across the range of $\lambda_{2}$ values by using the Bayesian Information Criterion (BIC). The model with the lowest BIC is selected as the final model, as it achieves the best trade-off between goodness of fit and model complexity, minimizing the risk of overfitting.
We choose BIC over other criteria based on empirical findings in Zhang et al. (2024), where pBIC was found to be a more conservative criterion. While a conservative selection method can be beneficial in some cases, detecting shifts in variance is inherently more challenging than detecting shifts in optimal values. A less conservative criterion like BIC is preferable here, as it increases sensitivity to shifts in variance while still penalizing overly complex models. Additionally, BIC offers a computational advantage over pBIC.